# The Angular Momentum Penrose Inequality

### v2

@@ Abstract
The proof introduces a four-stage Jang–conformal–AMO method: (1) solve an axisymmetric Jang equation with twist as a lower-order perturbation; (2) solve an angular-momentum-modified Lichnerowicz equation; (3) establish angular momentum conservation via de Rham cohomology; (4) apply the Dain–Reiris sub-extremality bound. The key innovation is the AM-Hawking mass $m_{H,J}(t):=\sqrt{m_{H}^{2}(t)+4\pi J^{2}/A(t)}$, which is monotonically non-decreasing along the $p$-harmonic flow and converges to $M_{\mathrm{ADM}}$.
As an application, we prove the Charged Penrose Inequality $M_{\mathrm{ADM}}\geq M_{\mathrm{irr}}+Q^{2}/(4M_{\mathrm{irr}})$ for non-rotating Einstein–Maxwell data.
Keywords: Penrose inequality $\cdot$ Angular momentum $\cdot$ Kerr spacetime $\cdot$ Marginally outer trapped surface $\cdot$ Cosmic censorship $\cdot$ Jang equation $\cdot$ Dominant energy condition
Mathematics Subject Classification (2020): Primary 83C57; Secondary 53C21, 83C05, 35J60, 58J05
@@ 1.1 Historical Context and Physical Motivation
The Penrose inequality, conjectured by Roger Penrose in 1973 [46], encapsulates a fundamental principle of black hole physics: the ADM mass of an asymptotically flat spacetime provides a lower bound determined by the area of its black hole horizons:
@@
However, the classical formulation (1) does not account for the angular momentum of the black hole. For rotating (Kerr) black holes, angular momentum plays a crucial role in determining the horizon structure and is a conserved quantity under Einstein evolution.
Executive Summary. This paper proves the Angular Momentum Penrose Inequality $M_{\mathrm{ADM}}\geq\sqrt{\frac{A}{16\pi}+\frac{4\pi J^{2}}{A}}$ for asymptotically flat, axisymmetric initial data with a stable MOTS, under the dominant energy condition and vacuum in the exterior region. Equality holds if and only if the data arises from a Kerr slice. The proof uses a four-stage method: (1) solve an axisymmetric Jang equation with twist perturbation; (2) apply an AM-Lichnerowicz conformal transformation; (3) establish angular momentum conservation and monotonicity of an “AM-Hawking mass” along a $p$-harmonic flow; (4) use the Dain–Reiris sub-extremality bound to control boundary terms. The core innovation is the functional $m_{H,J}(t):=\sqrt{m_{H}^{2}+4\pi J^{2}/A(t)}$, which interpolates between horizon and infinity while remaining monotone.
@@
A Kerr black hole with mass $M$ and angular momentum $J=aM$ is called sub-extremal if $|a|<M$, extremal if $|a|=M$, and super-extremal (or naked singularity) if $|a|>M$. Equivalently, in terms of the dimensionless spin $\chi:=a/M=J/M^{2}$: sub-extremal means $|\chi|<1$. For an axisymmetric MOTS with area $A$ and Komar angular momentum $J$, the sub-extremality condition is $A\geq 8\pi|J|$, which is equivalent to the existence of a Kerr solution with matching $(A,J)$. The sub-extremality factor appearing in monotonicity formulas is $(1-64\pi^{2}J^{2}/A^{2})=(1-(8\pi|J|/A)^{2})\geq 0$.
@@ 1.2 Main Result
    Axisymmetry: There exists a Killing field $\eta=\partial_{\phi}$ generating rotations, with $\eta\neq 0$ on $M\setminus\Gamma$ where $\Gamma$ denotes the rotation axis. (The axis $\Gamma=\{\eta=0\}$ is a 1-dimensional submanifold where the Killing field vanishes; the condition $\eta\neq 0$ on $M\setminus\Gamma$ ensures the orbits are circles corresponding to physical rotation.)
@@
with equality if and only if the initial data arises from a slice of the Kerr spacetime with parameters $(M,a=J/M)$.
@@
  • (H1) DEC: Used in Stage 1 to ensure $R_{\bar{g}}\geq 0$ via the Bray–Khuri identity, which through the AM-Lichnerowicz equation yields $R_{\tilde{g}}=\Lambda_{J}\phi^{-12}\geq 0$ and the mass chain $M_{\mathrm{ADM}}(\tilde{g})\leq M_{\mathrm{ADM}}(\bar{g})\leq M_{\mathrm{ADM}}(g)$ (Proposition E.1). Note: the bound $\phi\geq 1$ is not required—see Remark 5.7.
@@
Remark 1.4 (Extension to Marginally Stable MOTS).
Although Theorem 1.2 is stated for strictly stable MOTS ($\lambda_{1}(L_{\Sigma})>0$), the proof extends to the marginally stable case ($\lambda_{1}(L_{\Sigma})=0$) with minor modifications:
  1. (i)
    Jang equation: For marginally stable MOTS, the cylindrical decay rate becomes $\beta_{0}=2$ (from subleading spectral terms) rather than $\beta_{0}=2\sqrt{\lambda_{1}}$. The Jang solution still exists with logarithmic blow-up (see Lemma 5.5, Step 4).
  2. (ii)
    Lichnerowicz equation: The operator $8\Delta_{\Sigma}+R_{\Sigma}$ governing the conformal factor is distinct from the MOTS stability operator. By Lemma 5.5, $\lambda_{0}(-8\Delta_{\Sigma}+R_{\Sigma})>0$ for any stable MOTS $\Sigma\cong S^{2}$, regardless of whether $\lambda_{1}(L_{\Sigma})=0$.
  3. (iii)
    All subsequent stages: The AMO flow, angular momentum conservation, and monotonicity arguments depend only on $R_{\tilde{g}}\geq 0$ and the cylindrical end geometry, both of which remain valid for marginally stable MOTS.
Thus, the inequality (2) holds for all stable MOTS (including the marginally stable case). The strict stability hypothesis (H4) is imposed for clarity of exposition and because generic astrophysical black holes satisfy $\lambda_{1}>0$.
Corollary 1.5 (Angular Momentum Penrose Inequality for Stable MOTS).
The conclusion of Theorem 1.2 holds under the weaker hypothesis:
  • (H4’)
    Stable outermost MOTS: The principal eigenvalue satisfies $\lambda_{1}(L_{\Sigma})\geq 0$ (including the marginally stable case $\lambda_{1}=0$).
That is, for asymptotically flat, axisymmetric initial data satisfying (H1)–(H3) and (H4’), with outermost stable MOTS $\Sigma$ of area $A$ and Komar angular momentum $J$:
$M_{\mathrm{ADM}}\geq\sqrt{\frac{A}{16\pi}+\frac{4\pi J^{2}}{A}},$
@@
The result is presently restricted to axisymmetric data sets. The non-axisymmetric case remains a major open problem: without a Killing field, there is no canonical definition of quasi-local angular momentum, and the twist perturbation analysis does not apply. Dynamical horizons and the case of multiple black holes are discussed as open problems in Section 10.
@@
    Lower bound in terms of TT-tensor: If $\sigma^{TT}\neq 0$ (transverse-traceless part of $K$), then
    $\delta_{PI}\geq c_{0}\int_{M}|\sigma^{TT}|^{2}\,dV_{g}$
    for an explicit constant $c_{0}>0$ depending on the geometry.
@@
Part (i) follows from the rigidity analysis: $\delta_{PI}=0$ requires $\Lambda_{J}=\frac{1}{8}|\sigma^{TT}|^{2}=0$ identically (Section 9). The quantitative version comes from tracking the mass deficit through the Jang–conformal construction.
@@
    Metric and extrinsic curvature: $(g,K)\in C^{k,\alpha}_{\mathrm{loc}}(M)\times C^{k-1,\alpha}_{\mathrm{loc}}(M)$ for some $k\geq 3$ and $\alpha\in(0,1)$. This ensures:
    • Well-definedness of scalar curvature $R_{g}\in C^{k-2,\alpha}$;
@@
    • $C^{1,\alpha}$ regularity of $p$-harmonic potentials via Tolksdorf–Lieberman theory.
@@
    MOTS regularity: The outermost MOTS $\Sigma$ is a $C^{k,\alpha}$ embedded surface (automatic from elliptic regularity when $g\in C^{k,\alpha}$).
@@
    Minimal regularity: The proof can be extended to $C^{2}$ metrics using distributional techniques, but we state Theorem 1.2 for $C^{3,\alpha}$ data for clarity.
@@
Remark 1.9 (Key Analytical Inputs).
For readers interested in verifying the proof, we summarize the principal analytical ingredients:
  1. (i)
    Jang equation existence/regularity: The axisymmetric Jang equation with twist admits a solution with controlled logarithmic blow-up near the MOTS (Theorem 4.12). This relies on Han–Khuri [29, Theorem 1.1] and the twist perturbation estimate (Lemma 4.14): the twist term scales as $O(s)$ near the MOTS versus $O(s^{-1})$ for principal terms.
  2. (ii)
[…]
  3. (iii)
    AMO monotonicity: The $p$-harmonic functional monotonicity of Agostiniani–Mazzieri–Oronzio [1, Theorem 1.1] extends to the AM-Hawking mass when $R_{\tilde{g}}\geq 0$ and angular momentum is conserved (Theorem 6.22).
  4. (iv)
    Area–angular momentum bound: The Dain–Reiris inequality $A\geq 8\pi|J|$ for stable axisymmetric MOTS [22, Theorem 1] provides sub-extremality control (Theorem 7.1).
The proof is modular: each input is used in exactly one stage, allowing independent verification.
@@
  • $\alpha$ (in $C^{k,\alpha}$): The Hölder exponent, a regularity parameter $\alpha\in(0,1)$ appearing in function space definitions.
@@
These two uses of $\alpha$ appear in different contexts (regularity vs. differential forms) and should cause no confusion, but we emphasize the distinction here. When both appear nearby, we write $C^{k,\alpha}$ for regularity and $\alpha_{J}$ for the Komar form.
@@
For initial data $(M^{3},g,K)$, define the angular momentum source term $\Lambda_{J}$ as follows. Let $\sigma^{TT}$ denote the transverse-traceless part of the extrinsic curvature $K$ with respect to the York decomposition [54]:
@@
On the Jang manifold $(\bar{M},\bar{g})$ with $\bar{g}=g+df\otimes df$, define:
$\Lambda_{J}:=\frac{1}{8}|\sigma^{TT}|^{2}_{\bar{g}},$ (3)
where the norm is computed using the Jang metric $\bar{g}$:
$|\sigma^{TT}|^{2}_{\bar{g}}=\bar{g}^{ik}\bar{g}^{jl}\sigma^{TT}_{ij}\sigma^{TT}_{kl}.$
Key properties:
@@
    $\Lambda_{J}=0$ if and only if $\sigma^{TT}=0$ (no gravitational wave content);
@@
    For Kerr, $\sigma^{TT}=0$, so $\Lambda_{J}=0$ and the AM-Lichnerowicz equation reduces to $8\Delta_{\bar{g}}\phi+R_{\bar{g}}\phi=0$;
@@
    For rotating data with $J\neq 0$, generically $\Lambda_{J}>0$ away from the axis.
Physical interpretation: The term $\Lambda_{J}$ encodes the “gravitational radiation content” of the initial data. For stationary rotating black holes (Kerr), this content vanishes. For dynamical data (e.g., binary black hole mergers), $\Lambda_{J}>0$ represents gravitational wave energy not yet radiated to infinity.
@@
We therefore view vacuum as the minimal natural hypothesis for the angular momentum Penrose inequality in the present framework. The charged extension (§10.1) shows how specific matter models (electrovacuum) can be incorporated when their angular momentum contributions are computable.
@@
Mathematical necessity of vacuum for rotating black holes.
The vacuum hypothesis is not merely a technical convenience—it reflects a fundamental difference between rotating and non-rotating Penrose inequalities:
  • Non-rotating case ($J=0$): The Huisken–Ilmanen and Bray proofs require only DEC. Angular momentum plays no role, so matter contributions to the momentum constraint do not affect the argument.
  • Rotating case ($J\neq 0$): The proof requires tracking a conserved angular momentum along the flow. The conservation law $J(\Sigma_{t})=J(\Sigma_{0})$ holds if and only if $\nabla^{i}(K_{ij}\eta^{j})=0$, which requires $\boldsymbol{j}_{\phi}=0$ (vacuum). With matter satisfying only DEC, angular momentum generically flows between the black hole and the surrounding matter, making $J$ time-dependent along the foliation.
This distinction explains why the angular momentum Penrose inequality requires stronger hypotheses than the non-rotating version. Extending to non-vacuum matter would require either (a) a modified monotonicity formula that tracks $J$-variations, or (b) restricting to special matter models with $\boldsymbol{j}_{\phi}=0$ (e.g., co-rotating perfect fluids).

### v3

@@ Abstract
The proof introduces a four-stage Jang–conformal–AMO method: (1) solve an axisymmetric Jang equation with twist as a lower-order perturbation; (2) solve an angular-momentum-modified Lichnerowicz equation; (3) establish angular momentum conservation via Stokes’ theorem for closed 2-forms; (4) apply the Dain–Reiris sub-extremality bound. The key innovation is the AM-Hawking mass $m_{H,J}(t):=\sqrt{m_{H}^{2}(t)+4\pi J^{2}/A(t)}$, which is monotonically non-decreasing along the $p$-harmonic flow and converges to $M_{\mathrm{ADM}}$.
As a methodological application, we outline a new proof strategy for the Charged Penrose Inequality $M_{\mathrm{ADM}}\geq M_{\mathrm{irr}}+Q^{2}/(4M_{\mathrm{irr}})$ for non-rotating Einstein–Maxwell data, demonstrating the versatility of the Jang–conformal–AMO framework. This charged case is presented as a proof outline; complete rigorous proofs exist in the literature [37, 71].
Keywords: Penrose inequality, angular momentum, Kerr spacetime, marginally outer trapped surfaces, Jang equation, conformal method, positive mass theorem
Mathematics Subject Classification (2020): 83C57, 53C21, 35J60, 83C40
@@ 1.1 Historical Context and Physical Motivation
The Penrose inequality, conjectured by Roger Penrose in 1973 [48], encapsulates a fundamental principle of black hole physics: black holes cannot be “underweight” for their size. It relates the ADM mass of an asymptotically flat spacetime to the area of its black hole horizons:
@@
However, the classical formulation (1) does not account for the angular momentum of the black hole. For rotating (Kerr) black holes, angular momentum directly affects the horizon structure and is a conserved quantity under Einstein evolution.
@@
A Kerr black hole with mass $M$ and angular momentum $J=aM$ is called sub-extremal if $|a|<M$, extremal if $|a|=M$, and super-extremal (or naked singularity) if $|a|>M$. Equivalently, in terms of the dimensionless spin $\chi:=a/M=J/M^{2}$: sub-extremal means $|\chi|<1$. For an axisymmetric MOTS with area $A$ and Komar angular momentum $J$, the sub-extremality condition is $A\geq 8\pi|J|$, which is equivalent to the existence of a Kerr solution with matching $(A,J)$. The sub-extremality factor appearing in monotonicity formulas is:
$1-\frac{64\pi^{2}J^{2}}{A^{2}}=1-\left(\frac{8\pi|J|}{A}\right)^{2}.$
Key algebraic fact: This factor is non-negative if and only if $A\geq 8\pi|J|$. Indeed, $1-(8\pi|J|/A)^{2}\geq 0$ is equivalent to $(8\pi|J|/A)^{2}\leq 1$, i.e., $8\pi|J|\leq A$. At the extremal limit ($A=8\pi|J|$), the factor vanishes; for sub-extremal configurations ($A>8\pi|J|$), it is strictly positive. The Dain–Reiris inequality [22] ensures this factor is non-negative for all stable MOTS in axisymmetric data satisfying DEC.
@@ 1.2 Main Result
    Axisymmetry: There exists a Killing field $\eta=\partial_{\phi}$ generating rotations, with $\eta\neq 0$ on $M\setminus\Gamma$ where $\Gamma$ denotes the rotation axis; [^1] The axis $\Gamma=\{\eta=0\}$ is a 1-dimensional submanifold (possibly with multiple components) where the Killing field vanishes. The condition $\eta\neq 0$ on $M\setminus\Gamma$ ensures the orbits of $\eta$ are circles, corresponding to physical rotation about the axis.
@@
with equality if and only if the initial data arises from a slice of the Kerr spacetime with parameters $(M,a=J/M)$. The equality characterization is proven in Theorem 9.1 (Section 9) using the Mars–Simon tensor.
Remark 1.3 (Scope of the Result).
Theorem 1.2 establishes the Angular Momentum Penrose Inequality (2) for initial data satisfying axisymmetry and a vacuum exterior. The axisymmetry is required for defining the Komar angular momentum and for the orbit-space reduction of the Jang equation. The vacuum condition ensures $d^{\dagger}\alpha_{J}=0$, which is essential for angular momentum conservation (Theorem 6.14). We discuss these restrictions further in Remark 1.6 and Section 10.
Remark 1.4 (External Inputs).
The proof relies on several established results: the Jang equation existence theory of Han–Khuri [29], the area-angular momentum inequality of Dain–Reiris [22], the $p$-harmonic flow monotonicity of Agostiniani–Mazzieri–Oronzio [1], and the Mars–Simon characterization of Kerr [83, 84]. Our contribution is the synthesis of these methods and the introduction of the AM-Hawking mass functional.
@@
[…]
@@
The result is presently restricted to axisymmetric data sets with a vacuum exterior. These two constraints are fundamental to the proof:
  1. 1.
    Vacuum requirement: The result is strictly limited to data with $\mu=|\boldsymbol{j}|=0$ in the exterior region. This is necessary for the conservation of $J$ along the flow (Theorem 6.14). In the presence of matter, the Komar angular momentum would drift, and the inequality might require modification—see Remark 1.13 for details.
  2. 2.
    Axisymmetry requirement: The proof relies on the existence of the Killing field $\eta=\partial_{\phi}$, without which the Komar angular momentum $J$ is undefined. The non-axisymmetric case remains a major open problem: there is no canonical definition of quasi-local angular momentum, and the twist perturbation analysis does not apply.
Dynamical horizons and the case of multiple black holes are discussed as open problems in Section 10.
Physical situations excluded by the hypotheses:
  • Matter outside the horizon: Accretion disks, neutron star matter, or any stress-energy in the exterior region violates (H3). Angular momentum transport by matter would invalidate $J$-conservation.
  • Multiple disconnected horizons: The theorem addresses a single outermost MOTS. Binary black hole configurations require the “area additivity” conjecture (see Section 10).
  • Non-axisymmetric perturbations: Gravitational wave perturbations, tidally deformed horizons, or any configuration lacking a rotational Killing field violates (H2). The quasi-local angular momentum is then undefined.
  • Marginally stable MOTS ($\lambda_{1}=0$): The strictly stable condition (H4) excludes extremal or marginally stable horizons. The function space analysis requires $\lambda_{1}>0$ for Fredholm theory.
  • Dynamical horizons: Time-dependent horizons (e.g., during merger) are not covered; the theorem applies to a fixed initial data slice.
  • Charged rotating black holes (Kerr-Newman): The combined case $J\neq 0$ and $Q\neq 0$ is conjectured (Conjecture 10.36) but not proven.
@@
    Lower bound in terms of Kerr deviation: If $\mathcal{S}_{(g,K)}\neq 0$ (the Kerr deviation tensor from Definition 1.11), then
    $\delta_{PI}\geq c_{0}\int_{M}|\mathcal{S}_{(g,K)}|^{2}\,dV_{g}$
    for an explicit constant $c_{0}>0$ depending on the geometry. Note: this bound involves the Kerr deviation tensor, not the raw $\sigma^{TT}$, since Kerr slices themselves have $\sigma^{TT}\neq 0$.
@@
Part (i) follows from the rigidity analysis: $\delta_{PI}=0$ requires $\Lambda_{J}=\frac{1}{8}|\mathcal{S}_{(g,K)}|^{2}=0$ identically (Section 9). The quantitative version comes from tracking the mass deficit through the Jang–conformal construction.
@@
    Metric and extrinsic curvature: $(g,K)\in C^{k,\beta}_{\mathrm{loc}}(M)\times C^{k-1,\beta}_{\mathrm{loc}}(M)$ for some $k\geq 3$ and $\beta\in(0,1)$. This ensures:
    • Well-definedness of scalar curvature $R_{g}\in C^{k-2,\beta}$;
@@
    • $C^{1,\beta}$ regularity of $p$-harmonic potentials via Tolksdorf–Lieberman theory.
@@
    MOTS regularity: The outermost MOTS $\Sigma$ is a $C^{k,\beta}$ embedded surface (automatic from elliptic regularity when $g\in C^{k,\beta}$).
@@
    Minimal regularity: The proof can be extended to $C^{2}$ metrics using distributional techniques, but we state Theorem 1.2 for $C^{3,\beta}$ data for clarity.
@@
  • $\beta$ (in $C^{k,\beta}$): The Hölder exponent, a regularity parameter $\beta\in(0,1)$ appearing in function space definitions. We use $\beta$ rather than the traditional $\alpha$ to avoid any confusion with the Komar 1-form.
@@
This notational convention eliminates any potential ambiguity between regularity exponents and angular momentum-related quantities.
@@
For initial data $(M^{3},g,K)$, define the angular momentum source term $\Lambda_{J}$ as follows.
Preliminary: York decomposition. The extrinsic curvature $K$ admits the York decomposition [58]:
@@
Important clarification on Kerr geometry: Generic spacelike slices of the Kerr spacetime (e.g., Boyer–Lindquist $t=\mathrm{const}$ slices) are not conformally flat and possess non-trivial $\sigma^{TT}\neq 0$. This is in contrast to Bowen–York initial data, which is conformally flat by construction but does not represent exact Kerr slices. The condition $\sigma^{TT}=0$ characterizes conformally flat data, not Kerr data.
Definition of $\Lambda_{J}$ via Kerr deviation tensor. To correctly characterize the equality case, we define $\Lambda_{J}$ using the Kerr deviation tensor—a coordinate-independent object that vanishes if and only if the data is a Kerr slice. On the Jang manifold $(\bar{M},\bar{g})$ with $\bar{g}=g+df\otimes df$, define:
$\Lambda_{J}:=\frac{1}{8}|\mathcal{S}_{(g,K)}|^{2}_{\bar{g}},$ (3)
where $\mathcal{S}_{(g,K)}$ is the Kerr deviation tensor—a symmetric 2-tensor constructed intrinsically from $(g,K)$ that vanishes if and only if the initial data arises from a slice of Kerr spacetime.
Construction of the Kerr deviation tensor $\mathcal{S}_{(g,K)}$ (see Appendix G for complete details):
The construction uses the Killing Initial Data (KID) framework of Beig–Chruściel [85] and the Simon tensor characterization of Kerr [84, 86, 87]:
  1. (i)
    Electric and magnetic Weyl tensors: Define intrinsically from $(g,K)$:
    $E_{ij}$ $:=R_{ij}-\tfrac{1}{3}Rg_{ij}+(\mathrm{tr}K)K_{ij}-K_{ik}K^{k}{}_{j}$,
    $B_{ij}$ $:=\epsilon_{i}{}^{kl}\nabla_{k}K_{lj}.$
  2. (ii)
    Complex Weyl tensor: $\mathcal{W}_{ij}:=E_{ij}+iB_{ij}$.
  3. (iii)
    Reference Kerr Weyl tensor: For given $(M,J)$, the Weyl tensor $\mathcal{W}^{\mathrm{Kerr}}_{ij}(M,J)$ is determined by asymptotic matching (coordinate-independent via ADM frame).
  4. (iv)
    Kerr deviation: $\mathcal{S}_{(g,K),ij}:=\mathcal{W}_{ij}-\mathcal{W}^{\mathrm{Kerr}}_{ij}(M,J)$.
Why this is well-defined for non-stationary data: Even if $(g,K)$ does not arise from a stationary spacetime, the Weyl tensors $(E,B)$ are intrinsic to $(g,K)$. The comparison to Kerr is made via asymptotic matching using $(M,J)$, which is coordinate-independent. The Bianchi constraints propagate this comparison throughout $M$; see Lemma 5.5 (Section 5) for the complete rigorous construction and Appendix G for background on the Mars–Simon characterization.
Key properties (proven in Appendix G):
@@
    Characterization of Kerr (Theorem G.14): $\Lambda_{J}=0$ iff $\mathcal{S}_{(g,K)}=0$ iff the data is isometric to a Kerr slice;
@@
    For Kerr slices: $\Lambda_{J}=0$ by construction, even though $\sigma^{TT}\neq 0$ for generic Kerr slices;
@@
    For non-Kerr rotating data, generically $\Lambda_{J}>0$ away from the axis;
  5. (v)
    The tensor $\mathcal{S}_{(g,K)}$ encodes the “non-stationarity content” of the initial data.
Physical interpretation: The term $\Lambda_{J}$ measures the deviation of the initial data from Kerr geometry—it vanishes for any slice of Kerr (regardless of the slicing), and is positive for dynamical configurations. This is the correct characterization for the equality case: Kerr saturates the inequality precisely because $\Lambda_{J}=0$ for Kerr, not because $\sigma^{TT}=0$.
Remark 1.12 (Why $\sigma^{TT}$ alone is insufficient).
A common misconception is that $\sigma^{TT}=0$ characterizes Kerr. This is false:
  • Kerr slices have $\sigma^{TT}\neq 0$: Boyer–Lindquist slices of Kerr are not conformally flat. The induced 3-metric has non-trivial Cotton tensor, and the extrinsic curvature has genuine TT-content encoding frame-dragging.
  • Bowen–York data has $\sigma^{TT}=0$: Bowen–York initial data [89] is conformally flat with $\sigma^{TT}=0$, but it does not represent a Kerr slice—its evolution produces gravitational radiation.
The correct characterization uses the Mars–Simon tensor, which vanishes for Kerr (any slice) but is non-zero for Bowen–York and other non-Kerr configurations.
@@
We therefore view vacuum as the minimal natural hypothesis for the angular momentum Penrose inequality in the present framework. Without vacuum, the Komar angular momentum $J$ is not conserved along homologous surfaces, rendering the inequality $M\geq f(A,J)$ ill-posed: which value of $J$ (horizon vs. ADM vs. intermediate) should appear? The charged extension (§10.1) shows how specific matter models (electrovacuum) can be incorporated when their angular momentum contributions are computable.